OpenStudy (anonymous):

Consider the exponential function y=b^x, b>0. Determine the value for b for which the graph of y=b^x intersects the line y=x.

OpenStudy (aum):

y = b^x y = x If the graphs intersect then b^x = x Solve for x.

OpenStudy (aum):

I meant solve for b.

myininaya (myininaya):

The only way I can see how to solve for b such that b^x=x is to do it just by guessing... There is one value b such that b^x=x from what I can tell.

myininaya (myininaya):

And I mean there is one value b such that b^x=x for some real value x

OpenStudy (anonymous):

what does it mean that it must intersect the line y=x?

myininaya (myininaya):

that one curve must go through the other curve is what intersects means

myininaya (myininaya):

you have heard of intersections in highways

myininaya (myininaya):

 look at that exponential equation it doesn't touch the y=x line

myininaya (myininaya):

There should be a b that you can think of so that y=b^x is not an exponential equation

OpenStudy (aum):

It will if 0 < b < 1

OpenStudy (aum):

exponential decay function instead of a growth function.

OpenStudy (aum):